OpenStudy (anonymous):

Three disjoint subsets are to be formed from a collection of 30 items. The first is to have 10 elements, the second is to have 9 elements, and the third is to have 11 elements. In how many ways can this be done?

OpenStudy (anonymous):

somehow my answer doesnt match up the answer from the book, strange

OpenStudy (anonymous):

\[\left(\begin{matrix}30 \\ 10\end{matrix}\right), \left(\begin{matrix}30 \\ 9\end{matrix}\right),\left(\begin{matrix}30 \\ 11\end{matrix}\right)\]

OpenStudy (anonymous):

I got 989779465 but the book answer is 5046360719400

OpenStudy (amistre64):

does order matter?

OpenStudy (amistre64):

i think 30C10, then after 10 are chosen its 20C9, etc

OpenStudy (amistre64):

\[\binom{n}{a}\binom{n-a}{b}\binom{n-a-b}{c}...\]

OpenStudy (anonymous):

k thank you for the clarification

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

Three disjoint subsets are to be formed from a collection of 30 items. The first is to have 11 elements, the second is to have 9 elements, and the third is to have 10 elements. In how many ways can this be done? that will give the same answer?